The typing speeds for secretaries are normally distributed. If a random sample of secretaries is taken and the confidence interval is (43.25,46.75), what is the error bound?
Answer #1

Confidence itnervals (43.25, 46.75)

To calculate the sample mean: \bar{X}= \frac{X_2+X_1}{2}=\frac{43.25+46.75}{2}=45

To calculate error bound: EBM= \frac{X_2-X_1}{2}= \frac{46.75-43.25}{2}=1.75